OpenStudy (anonymous):

Fan and Medal 12 - x = 7 x = ? Show Your Steps Please

OpenStudy (midhun.madhu1987):

Subtract 12 from both sides...

OpenStudy (ahsome):

\[12-x=7\] \[-x=7-12\] \[-x=-5\] \[x=5\]

OpenStudy (anonymous):

from 12 and 7 right

OpenStudy (ahsome):

Thats it :) Just say if you don't understand

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i like how you do yours it is less complicated

OpenStudy (anonymous):

but one question though

OpenStudy (anonymous):

will the 2 numbers always be subtracted

OpenStudy (ahsome):

Only if you move a positive number, @CJBALLER5323. In this case, since we moved positive 12, it became \(-12\). If we moved \(-12\), it would be add 12. Got that?